when equal volumes of liquid AandB are mixed,a solution with a density of 1.2g/mL is obtained.when equal masses of these liquids are mixed the density of the mixture is found to be o.9g/mL.find the density of each kliquid.
1
Expert's answer
2012-12-03T09:43:45-0500
d = (d1+d2)/2

d1 = m1/v1 v1 = m1/d1
d2 = m2/v2 v2 = m2/d2

when v1 = v2, d = (d1+d2)/2

when m1 = m2 d = 2m1/(m1/d1+m2/d2) = 2m1/((m1d2+m2d1)/d1d2) = 2d1d2/d1+d2

1.2 = (d1+d2)/2
0.9 = 2d1d2/d1+d2

0.9 = 2d1d2/d1+d2
2.4 = d1+d2

0.9 = 2d1d2/d1+d2
d1 = d2-2.4

0.9 = 2( d2-2.4 )d2/ (d2-2.4 )+d2

d1 = 1.8
d2 = 0.6